Consider the following functions.
f(x) =
9 − 6x
, g(x) =
x + 7

(a) Find (f + g)(x), (f − g)(x), (fg)(x), and (f/g)(x).
(f + g)(x) =
(f − g)(x) =
(fg)(x) =
(f/g)(x) =

Answer by sheldonbbtrocks(115)

f+g(x) means add the functions:
9-6x+x+7=-5x+16
(f + g)(x) = [-5x+16]

(f-g)(x) means subtract g from f:
(9-6x)-(x+7)=-7x+2
(f − g)(x) = [-7x+2]
(fg)(x) means multiply the functions:
(9-6x)(x+7)=9x+63-6x^2-42=
(fg)(x) = [-6x^2+3x+21]

(f/g)(x) means divide g by f:
(9-6x)/(x+7)
(f/g)(x) =[(9-6x)/(x+7)]
Sorry this is a little messy, I'll rewrite the answers here:
(f + g)(x) = [-5x+16]
(f - g)(x) = [-7x+2]
(f * g)(x) = [-6x^2+3x+21]
(f / g)(x)= [(9-6x)/(x+7)]
Hope this helped!